a rectangle's length is equal to eight times its width.if the area of the rectangle is 392 square inches, what is the width of the rectangle?
Answer by checkley77(12844)

L=8W
A=LW
392=8W*W
392=8W^2
8W^2=392
W^2=392/8
W^2=49
W=SQRT49
W=7 ANS.
L=8*7=56 ANS.
PROOF:
392=7*56
392=292